If an egg carton has 12 eggs and I fill 3 5/12 of the cartons with eggs how many eggs do I have altogether


1
Expert's answer
2021-02-24T06:40:53-0500

"12*3\frac{5}{12}=12*\frac{41}{12}=41\; eggs"